OpenStudy (anonymous):

Can anyone help? Dan purchases DVD's from a website. Each DVD costs $11, and the shipping and handling fees are $6.95. Dan is charged a total of $50.95. How many DVD's did he purchase??

OpenStudy (anonymous):

well first off you need to subtract the 6.95 from your total

OpenStudy (texaschic101):

11d + 6.95 = 50.95

OpenStudy (anonymous):

Thank you guys so much!

OpenStudy (anonymous):

11x=50.95-6.95

OpenStudy (anonymous):

No prob

OpenStudy (anonymous):

Then combine like terms. 11x=44 Solve for x by dividing 11 on both sides. \[x=44\div11\]

OpenStudy (anonymous):

So x is the number of dvd's he purchased. :D

OpenStudy (anonymous):

You can use the equation y=mx+b y=total money x=11 dollars m=number of videos b= handling fee therefore 50.95=m(11)+6.95 then solve for m

OpenStudy (texaschic101):

I just set the equation up different...but it will be the same answer

OpenStudy (anonymous):

Yep just used a different equation :)